OpenStudy (anonymous):

Which property of potential energy distinguishes it from kinetic energy? A. mass B. position C. force D. work and power

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

i have to finish this up before i can eat dinner so :P

OpenStudy (anonymous):

i think it is force

jagr2713 (jagr2713):

correct

OpenStudy (anonymous):

Compare the variables in the kinetic energy equation to the variables in the potential energy equations. Gravitational Potential Energy\[=mass \times gravitational\ constant \times displacement\]Elastic Potential Energy \[= \frac{ 1 }{ 2 }\times spring\ constant \times (displacement)^{2} \]Kinetic Energy \[= \frac{ 1 }{ 2 }\times mass \times (velocity)^{2}\]

OpenStudy (anonymous):

Not sure here, but I think b, position should be correct. The reason being is because: Can't be a, since you need mass to calculate both type of energy Can't be c, because both type of energy requires a force (kinetic energy is obvious because it's moving), but you also need a force to force the object to hold/store potential energy (either in the form of normal force against gravity, or electrostatic energy if you want to force a particle to stay still between charged plates). can't be d, because an object with a constant kinetic energy will have 0 work done on it, so in that sense it's same with potential energy. b makes sense for me since if you vary the position, you change the potential energy but the kinetic energy would be the same